# Conformal Prediction for Time-series Forecasting with Change Points

### v1

@@ 2 Related Work
Conformal prediction (CP) (vovk2005algorithmic, ) has emerged as a leading framework for UQ due to its algorithmic simplicity, broad applicability, and finite-sample coverage guarantees; see angelopoulos2021gentle for a comprehensive introduction. For time-series forecasting, some works focus on joint coverage over fixed-length horizons (Schaar2021conformaltime, ; sun2023copula, ; zhou2024heterogeneous, ), while others assume stable underlying dynamics and adapt only to shifts in residual distributions (xu2021conformal, ) . tibshirani2019conformal uses known covariate shifts (by propensity weighting) to achieve theoretical coverage, but can fail in settings where the shift mechanism is unknown, which is typical for time series forecasting. Conformal prediction has also been extended for change-point detection in time series via conformal martingales (vovk2021retrain, ; volkhonskiy2017inductive, ).
@@ 4 Conformal Prediction for Time series with Change points (CPTC)
We consider the online conformal prediction setting where we have a sequence of data points from $\mathcal{X}\times\mathcal{Y}$: $(x_{1},y_{1}),$ $(x_{2},y_{2}),\ldots,(x_{T},y_{T})$. We have a trained state predictor for $\hat{p}(z_{t}|x_{t})$ , a state-conditioned forecaster $\hat{f}:{\mathcal{X}}\times{\mathcal{Z}}\to{\mathcal{Y}}$ (a probabilistic SDS model can serve both purposes), and are given a desired confidence level $1-\alpha$. At each time step $t$, we have seen $(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{t-1},y_{t-1})$ and the next input feature vector $x_{t}$. We want to produce an adaptive, time-dependent prediction interval based on input $\Gamma_{t}(x_{t})\subseteq\mathcal{Y}$ for the unseen true target $y_{t}$, which we will learn at the next time step. Let $z_{t}\in{\mathcal{Z}}$ be the true (latent switching) state at time $t$. Our goal is to produce prediction intervals with correct empirical coverage, i.e.
@@ 4.1 Algorithm
Input: nonconformity score function AA, probabilistic state model $\hat{p}(z_{t}|x_{t})$ with prior $\hat{p}(z_{0})$, forecaster model $\hat{f}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$, confidence level $1-\alpha$, learning rate $\gamma$
@@ 3   Initialize ${\mathcal{S}}_{z}\leftarrow\{\}$, $\alpha_{z,0}\leftarrow\alpha$
7    ${\mathcal{S}}_{z_{t}}\leftarrow{\mathcal{S}}_{z_{t}}\cup\{A(x_{t},y_{t})\}$, where $z_{t}\sim\hat{p}(z_{t}=z|x_{t})$; // warm start scores
@@ 17 $\Gamma_{t}(x_{t})\leftarrow$ Aggregate the $\Gamma_{z,t}(x_{t})$s by Eqn 10 or Eqn 11;
18   Sample state $\hat{z}_{t}\sim\hat{p}(z_{t}|x_{t})$; // state-specific coverage target tracking
@@
$\sum_{z\in{\mathcal{Z}}}P(y_{t}\in\Gamma_{t}(x_{t})|z_{t}=z)\;P(z_{t}=z|x_{t})\geq 1-\alpha$ (7)
@@
For each switching state $z\in Z$, we initialize (1) a mode-specific set of nonconformity scores ${\mathcal{S}}_{z}$ and (2) the mode-specific confidence level $\alpha_{z,0}=\alpha$. If there are $w$ steps of warm-start data $\{(x_{t},y_{t})\}_{t=-w}^{0}$, we initialize the sore sets ${\mathcal{S}}_{z}$ by sampling $z_{t}\sim\hat{p}(z_{t}=z|x_{t})$ for $t=-w,\ldots,0$, and inserting $A(x_{t},y_{t})$ into ${\mathcal{S}}_{z_{t}}$. Depending on the size of the warm-start window and desired properties, practitioners can adjust the warm-starting approach. For example, using the entire warm-start window ${\mathcal{S}}_{z}=\{A(x_{t},y_{t})\}_{t=-w}^{0}$ for all $z\in{\mathcal{Z}}$ will result in better stability but less adaptability to modes. The approach does not change the algorithm’s coverage guarantees.
@@
At every time step $t\in[1,T]$, CPTC creates state-specific conformal prediction intervals for all states with nonzero probability. This allows us to generate predictions tailored to the current anticipated dynamics. Specifically, for every state $z\in{\mathcal{Z}}$ where $\hat{p}(z|x_{t})>0$, we first obtain the point prediction through forecaster model $\hat{f}(x_{t},z)$, and then the state-specific prediction interval $\Gamma_{z,t}(x_{t})$ by calibrating on nonconformity scores set ${\mathcal{S}}_{z}$ to the adaptive confidence level of $1-\alpha_{z,t}$.
@@
Here $\gamma>0$ is the step size hyperparameter, and $\hat{z}_{t}\in{\mathcal{Z}},\;\hat{z}_{t}\sim\hat{p}(z_{t}=z|x_{t})$ is the state at time $t$ sampled to be updated. The update rule increases $\alpha_{\hat{z}_{t},t}$ if coverage is too conservative and decreases it otherwise, ensuring that for each $z\in{\mathcal{Z}}$, $\alpha_{z,t}$ converges to a value that maintains long-term coverage.
@@ 4.2 Theoretical analysis
In real-world applications, the state predictor $\hat{p}(z_{t}|x_{t})$ that plays a large role in our algorithm is often imperfect. The CPTC algorithm accommodates these scenarios by tracking $\alpha^{*}_{z}$ for each predicted mode $z$ with adaptive online updates. Theorem 4.3 quantifies the effect of imperfect state prediction by showing that the miscoverage rate is bounded by the product of the misclassification rate and state-specific deviations.
@@ 5 Experiments
Table 1: Performance on synthetic and real-world datasets with target confidence $1-\alpha=0.9$ (for horizon $T=200$ for synthetic datasets, $400$ for electricity and traffic, and $60$ for bee, mean $\pm$ standard deviation of the test samples). Methods that are invalid (coverage below $90\%$) are grayed out. Our method achieves a high level of calibration (coverage is close to $90\%$) consistently.
@@ 5.1 Analysis of Coverage and Sharpness Results
Results for uncertainty quantification in table 1 show that we achieve significantly better validity on all datasets compared to baselines. RED-SDS’s intervals fail due to lack of calibration, a common pitfall of probabilistic methods (shown also in xu2023sequential ). CP’s under-coverage shows that the data are not exchangeable. In the pretense of change points, ACI’s coverage fluctuates dramatically and did not converging within the horizon (more visualizations in Appendix B.5). For SPCI and HopCPT, which also guarantees asymptotic validity, the invalidity is likely because they are designed for large datasets ($T\geq 10,000$ for HopCPT), and did not have enough data to react or learn useful residual patterns for the relatively short prediction horizon of our datasets ($T=200$ for synthetic datasets, 400 for electricity and traffic, and 60 for bees). Figure 3 provides a qualitative example of how the prediction intervals compare across different methods. Our method (purple shaded regions) does not over-cover during nonvolatile hours, nor under-cover during busy ones, as does ACI.
@@ 5.2 Ablation Studies
Table 2: Coverage of CPTC for using ground truth (GT) labels and with various levels of injected noise, on synthetic data with $T=500$. We can see that coverage performance does not change significantly with error in state prediction.
@@ B.1 Hyperparameters for the base forecasting model
We train our RED-SDS on the synthetic datasets, and use the model checkpoints provided by the authors for the real-world datasets. The model architecture consists of a discrete switching component with $K\in\{2,3\}$ categories and a continuous state space with dimensionality $d_{x}\in\{2,4\}$. For training, we employ the ELBOv2 objective with a learning rate $\eta\in[5\times 10^{-3},7\times 10^{-3}]$, warmup steps of 1000, and gradient clipping at 10.0. The model uses a batch size $B\in\{32,50\}$ and is trained for $T\in\{20,000,30,000\}$ steps. The continuous transition and emission models are parameterized by nonlinear MLPs with hidden dimensions $h=32$, while the inference network uses either a bidirectional RNN or transformer with embedding dimensions $d_{e}=4$. For real-world datasets, we apply target transformation and Jacobian correction, while synthetic datasets use raw observations. The model’s capacity is controlled through weight decay $\lambda=10^{-5}$ and MLP hidden dimensions $h\in\{8,32,64\}$ depending on the component. For forecasting, the model is trained to forecast a window of $t=50$ time steps for synthetic datasets (bouncing ball) and a windows of for real-world datasets (electricity, traffic) as specified by their respective metadata. The model generates $N=100$ Monte Carlo samples for each forecast, with deterministic latent states ($z_{t}$) and stochastic continuous states ($x_{t}$) to capture both the switching dynamics and the inherent uncertainty in the predictions. We segment all datasets by a 70/10/20 train/validation/test split. Conforaml prediction results are reported on the test set only.
@@ B.4 Additional Experiments
Coverage (union) 91.03 90.44 93.35 94.76 92.62 93.43
Coverage (min) 91.03 90.35 91.50 90.87 89.82 90.28
Width (union) 2.13 4.92 13.68 8.34 166.88 1.15
Width (min) 2.13 4.85 12.43 7.95 153.57 1.08

### v2

@@ 1 Introduction
Our code is open source at https://github.com/huiwenn/CPTC. Our contributions are:
@@ 2 Related Work
Conformal prediction (CP) (vovk2005algorithmic, ) has emerged as a leading framework for UQ due to its algorithmic simplicity, broad applicability, and finite-sample coverage guarantees; see angelopoulos2021gentle for a comprehensive introduction. For time-series forecasting, some works focus on joint coverage over fixed-length horizons (Schaar2021conformaltime, ; sun2023copula, ; zhou2024heterogeneous, ); others assume stable underlying dynamics and adapt only to shifts in residual distributions (xu2021conformal, ) . tibshirani2019conformal uses known covariate shifts (by propensity weighting) to achieve theoretical coverage, but can fail in settings where the shift mechanism is unknown, which is typical for time series forecasting. hajihashemi2024multi introduces a multi-model ensemble CP framework, optimizing model selection within a pool of experts for combined coverage. CP has also been extended for change-point detection in time series via conformal martingales (vovk2021retrain, ; volkhonskiy2017inductive, ).
@@ 4 Conformal Prediction for Time series with Change points (CPTC)
We consider the online conformal prediction setting where we have a sequence of data points from $\mathcal{X}\times\mathcal{Y}$: $(x_{1},y_{1}),$ $(x_{2},y_{2}),\ldots,(x_{T},y_{T})$. We have a trained state predictor for $\hat{p}(z_{t}|x_{0:t})$ , a state-conditioned forecaster $\hat{f}:{\mathcal{X}}\times{\mathcal{Z}}\to{\mathcal{Y}}$ (a probabilistic SDS model can serve both purposes), and are given a desired confidence level $1-\alpha$. At each time step $t$, we have seen $(x_{1},y_{1}),(x_{2},y_{2}),\ldots,(x_{t-1},y_{t-1})$ and the next input feature vector $x_{t}$. We want to produce an adaptive, time-dependent prediction interval based on input $\Gamma_{t}(x_{t})\subseteq\mathcal{Y}$ for the unseen true target $y_{t}$, which we will learn at the next time step. Let $z_{t}\in{\mathcal{Z}}$ be the true (latent switching) state at time $t$. Our goal is to produce prediction intervals with correct empirical coverage, i.e.
@@ 4.1 Algorithm
Input: nonconformity score function AA, probabilistic state model $\hat{p}(z_{t}|x_{0:t})$ with prior $\hat{p}(z_{0})$, forecaster model $\hat{f}:\mathcal{X}\times\mathcal{Z}\rightarrow\mathcal{Y}$, confidence level $1-\alpha$, learning rate $\gamma$
@@ 3   Initialize ${\mathcal{S}}_{z}\leftarrow\{\}$, $\alpha_{z,0}\leftarrow\alpha$
7    ${\mathcal{S}}_{z_{t}}\leftarrow{\mathcal{S}}_{z_{t}}\cup\{A(x_{t},y_{t})\}$, where $z_{t}\sim\hat{p}(z_{t}=z|x_{0:t})$; // warm start scores
@@ 17 $\Gamma_{t}(x_{t})\leftarrow$ Aggregate the $\Gamma_{z,t}(x_{t})$s by Eqn 10 or Eqn 11;
18   Sample state $\hat{z}_{t}\sim\hat{p}(z_{t}|x_{0:t})$; // state-specific coverage target tracking
@@
$\sum_{z\in{\mathcal{Z}}}P(y_{t}\in\Gamma_{t}(x_{t})|z_{t}=z)\;P(z_{t}=z|x_{0:t})\geq 1-\alpha$ (7)
@@
For each switching state $z\in Z$, we initialize (1) a mode-specific set of nonconformity scores ${\mathcal{S}}_{z}$ and (2) the mode-specific confidence level $\alpha_{z,0}=\alpha$. If there are $w$ steps of warm-start data $\{(x_{t},y_{t})\}_{t=-w}^{0}$, we initialize the sore sets ${\mathcal{S}}_{z}$ by sampling $z_{t}\sim\hat{p}(z_{t}=z|x_{0:t})$ for $t=-w,\ldots,0$, and inserting $A(x_{t},y_{t})$ into ${\mathcal{S}}_{z_{t}}$. Depending on the size of the warm-start window and desired properties, practitioners can adjust the warm-starting approach. For example, using the entire warm-start window ${\mathcal{S}}_{z}=\{A(x_{t},y_{t})\}_{t=-w}^{0}$ for all $z\in{\mathcal{Z}}$ will result in better stability but less adaptability to modes. The approach does not change the algorithm’s coverage guarantees.
@@
At every time step $t\in[1,T]$, CPTC creates state-specific conformal prediction intervals for all states with nonzero probability. This allows us to generate predictions tailored to the current anticipated dynamics. Specifically, for every state $z\in{\mathcal{Z}}$ where $\hat{p}(z|x_{0:t})>0$, we first obtain the point prediction through forecaster model $\hat{f}(x_{t},z)$, and then the state-specific prediction interval $\Gamma_{z,t}(x_{t})$ by calibrating on nonconformity scores set ${\mathcal{S}}_{z}$ to the adaptive confidence level of $1-\alpha_{z,t}$.
@@
Here $\gamma>0$ is the step size hyperparameter, and $\hat{z}_{t}\in{\mathcal{Z}},\;\hat{z}_{t}\sim\hat{p}(z_{t}=z|x_{0:t})$ is the state at time $t$ sampled to be updated. The update rule increases $\alpha_{\hat{z}_{t},t}$ if coverage is too conservative and decreases it otherwise, ensuring that for each $z\in{\mathcal{Z}}$, $\alpha_{z,t}$ converges to a value that maintains long-term coverage.
@@ 4.2 Theoretical analysis
In real-world applications, the state predictor $\hat{p}(z_{t}|x_{0:t})$ that plays a large role in our algorithm is often imperfect. The CPTC algorithm accommodates these scenarios by tracking $\alpha^{*}_{z}$ for each predicted mode $z$ with adaptive online updates. Theorem 4.3 quantifies the effect of imperfect state prediction by showing that the miscoverage rate is bounded by the product of the misclassification rate and state-specific deviations.
@@ 5 Experiments
Table 1: Performance on synthetic and real-world datasets with target confidence $1-\alpha=0.9$ (for horizon $T=200$ for synthetic datasets, $300$ for electricity and traffic, and $60$ for bee, mean $\pm$ standard deviation of the test samples). Methods that are invalid (coverage below $90\%$) are grayed out. Our method achieves a high level of calibration (coverage is close to $90\%$) consistently.
@@ 5.1 Analysis of Coverage and Sharpness Results
Results for uncertainty quantification in table 4 show that we achieve significantly better validity on all datasets compared to baselines. RED-SDS’s intervals fail due to lack of calibration, a common pitfall of probabilistic methods (shown also in xu2023sequential ). CP’s under-coverage shows that the data are not exchangeable. In the pretense of change points, ACI’s coverage fluctuates dramatically and did not converging within the horizon (more visualizations in Appendix B.5). For SPCI and HopCPT, which also guarantees asymptotic validity, the invalidity is likely because they are designed for large datasets ($T\geq 10,000$ for HopCPT), and did not have enough data to react or learn useful residual patterns for the relatively short prediction horizon of our datasets ($T=200$ for synthetic datasets, 300 for electricity and traffic, and 60 for bees). Figure 3 provides a qualitative example of how the prediction intervals compare across different methods. Our method (purple shaded regions) does not over-cover during nonvolatile hours, nor under-cover during busy ones, as does ACI.
@@ 5.2 Ablation Studies
Table 2: Coverage of CPTC for using ground truth (GT) labels and with various levels of injected noise, on synthetic data with $T=200$. We can see that coverage performance does not change significantly with error in state prediction.
@@ B.1 Hyperparameters for the base forecasting model
We train our RED-SDS on the synthetic datasets, and use the model checkpoints provided by the authors for the real-world datasets. The model architecture consists of a discrete switching component with $K\in\{2,3\}$ categories and a continuous state space with dimensionality $d_{x}\in\{2,4\}$. For training, we employ the ELBOv2 objective with a learning rate $\eta\in[5\times 10^{-3},7\times 10^{-3}]$, warmup steps of 1000, and gradient clipping at 10.0. The model uses a batch size $B\in\{32,50\}$ and is trained for $T\in\{20,000,30,000\}$ steps. The continuous transition and emission models are parameterized by nonlinear MLPs with hidden dimensions $h=32$, while the inference network uses either a bidirectional RNN or transformer with embedding dimensions $d_{e}=4$. For real-world datasets, we apply target transformation and Jacobian correction, while synthetic datasets use raw observations. The model’s capacity is controlled through weight decay $\lambda=10^{-5}$ and MLP hidden dimensions $h\in\{8,32,64\}$ depending on the component. For forecasting, the model is trained to forecast a window of $t=50$ time steps for synthetic datasets (bouncing ball) and a windows of for real-world datasets (electricity, traffic) as specified by their respective metadata. The model generates $N=100$ $(x_{t},y_{t},z_{t})_{t=0}^{k}$ triplet trajectories from the trained model by Monte Carlo sampling, and use the samples to calculate prediction quantiles for table 1, in accordance to the original paper.s. We segment all datasets by a 70/10/20 train/validation/test split. Conforaml prediction results are reported on the test set only.
@@ B.4 Additional Experiments
Additional Baselines
We add comparison to three more baselines that handles distribution shift in time series data. They are DtACI [5], ECI [55], and L-ARC [61].
We would like to emphasize that these baselines are reactive. The question we try to address with CPTC is - what happens when we can anticipate distribution shift? CPTC is designed such that it greatly improves coverage compared to purely reactive methods, while still maintaining the ability to adapt to unknown shifts. To our knowledge, there are no other models that explore the same problem; CPTC provides practitioners with a simple and sound solution, when they have the ability (such as via a state space model) to predict shifts.
Table 4: Performance on synthetic and real-world datasets with target confidence $1-\alpha=0.9$ (for horizon $T=200$ for synthetic datasets, $300$ for electricity and traffic, and $60$ for bee, mean $\pm$ standard deviation of the test samples). Methods that are invalid (coverage below $90\%$) are grayed out. Our method achieves a high level of calibration (coverage is close to $90\%$) consistently.
DtACI ECI-cutoff L-ARC Ours
Bouncing
Ball obs.
Cov 82.97 $\pm$ 1.35 79.86 $\pm$ 1.42 75.12 $\pm$ 9.25 92.06 $\pm$ 1.93
Width 3.21 $\pm$ 0.80 3.02 $\pm$ 0.72 1.89 $\pm$ 0.76 10.70 $\pm$ 3.01
Bouncing
Ball dyn.
Cov 79.63 $\pm$ 1.10 80.91 $\pm$ 1.28 78.63 $\pm$ 6.21 90.62 $\pm$ 3.85
Width 3.51 $\pm$ 2.20 2.32 $\pm$ 0.19 2.44 $\pm$ 0.23 10.73 $\pm$ 3.67
3-Mode
System
Cov 88.82 $\pm$ 0.95 83.67 $\pm$ 0.89 92.58 $\pm$ 2.77 90.43 $\pm$ 1.92
Width 12.55 $\pm$ 7.52 2.39 $\pm$ 0.48 12.31 $\pm$ 2.85 4.80 $\pm$ 2.64
Traffic
Cov 89.14 $\pm$ 0.40 86.12 $\pm$ 0.50 90.05 $\pm$ 6.24 92.38 $\pm$ 1.24
Width 10.35 $\pm$ 9.85 7.02 $\pm$ 9.72 8.39 $\pm$ 10.13 15.83 $\pm$ 5.96
Electricity
Cov 88.35 $\pm$ 0.38 89.41 $\pm$ 0.47 86.42 $\pm$ 4.96 91.19 $\pm$ 1.22
Width 189.51 $\pm$ 187.25 138.72 $\pm$ 83.30 157.82 $\pm$ 95.02 249.46 $\pm$ 1365.46
Dancing
Bees
Cov 78.91 $\pm$ 3.02 80.12 $\pm$ 2.85 75.92 $\pm$ 14.21 95.71 $\pm$ 4.95
Width 0.03 $\pm$ 0.01 0.08 $\pm$ 0.01 0.28 $\pm$ 0.15 0.33 $\pm$ 0.14
@@
Coverage (union) 90.53 90.44 93.35 92.76 92.62 93.43
Coverage (min) 90.03 90.32 91.50 91.87 89.82 90.28
Width (union) 3.73 1.72 2.45 7.94 166.88 0.79
Width (min) 3.62 1.65 2.41 7.95 153.57 0.77
Discussion on Misspecified Regimes.
We designed the number of latent states to be a hyperparameter given by the user. Our method ensures that the validity guarantee still holds even if the state cardinality is incorrect (theorem 4.2), as incorrect state cardinality can be formulated as a case of incorrect classification. Determining state space cardinality is a well-studied problem (see for example [14]), and we consider it to be beyond the scope of our paper.
Regarding the effect on performance, our ablation studies (Section 5.2, Tables 2 and 3) demonstrate that while the validity (coverage) of CPTC is robust to errors in state prediction, the sharpness (width of prediction intervals) can be affected. Table 3 explicitly shows that ”the more accurate the state prediction is, the sharper the intervals.” This implies that selecting a more appropriate number of regimes that better captures the true underlying dynamics would lead to narrower and more efficient prediction intervals.
@@ B.5 Qualitative Results
Appendix C Runtime analysis
We provide a runtime analysis comparing the aggregation methods in Equations 10 and 11 and with baselines.
Notation
  • $T$: Total timesteps, $K$: Number of states, $M$: State prediction model inference complexity
  • Score sets $\mathcal{S}_{z}$ maintained in sorted order (quantile computation in $\mathcal{O}(1)$, insertion $\mathcal{O}(\log T)$)
Complexity Per Timestep
  1. 1.
    Computing quantile $Q^{1-\alpha_{t}}(\mathcal{S}_{z}^{(l)})$ for each state, totaling $\mathcal{O}(K)$.
  2. 2.
    State Prediction: computing $\hat{p}(z_{t}|x_{t})$ and sampling a state, totaling $\mathcal{O}(M)$.
  3. 3.
    Aggregation
    • Eq. 10, Weighted Level-Set: $\mathcal{O}(K\cdot\lceil 1/\delta\rceil^{d})$ where $\delta$ is discretization resolution
    • Eq. 11, Union: $\mathcal{O}(K\log K)$ for sorting states by probability
  4. 4.
    Updates: $\mathcal{O}(\log t)$ for inserting new score into sorted set.
Therefore, the total complexity of CPTC with horizon $T$ is:
$\mathcal{O}(T\cdot(M+K\cdot\lceil\frac{1}{\delta}\rceil^{d})+T\log T)$ if we use the Weighted Level-Set as aggregation in Eq. 10, and
$\mathcal{O}(T\cdot(M+K\log K)+T\log T)$ if we use the more efficient Union aggregation in Eq. 11.
In practice, both $M$ (inference cost) and $K$ (number of discrete states) are small. For applications where $K=\mathcal{O}(1)$ and $M=\mathcal{O}(1)$, CPTC achieves near-linear scaling $\mathcal{O}(T\log T)$, the same as ACI (see table below).
If we use the more precise weighted level set aggregation method, the runtime largely depends on the discretization resolution $\delta$, output dimension $d$, and the number of states $K$. We remark the level set approach scales poorly if the resolution, dimension, or number of states are high.
Method Time Complexity Remarks
ACI $\mathcal{O}(T\log T)$ Single score set, no state awareness
SPCI / HopCPT $\mathcal{O}(T^{2}\log T)$ Re-train regression model at every time step
CPTC (Union) $\mathcal{O}(T\log T)$ State-aware, pre-trained predictor